the force acting required to tow a boat at constant velocity is directky proportional to speed...if a speed 4 km/hr,corresponding power is 7.5kw....how much power is neede for speed=12km/hr?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shivanshi

given

F= α V

power P = FV

            =α V2

      7.5  =α * 16

        α   =7.5/16

so at 12 km/hr

   P = α * 144

       =7.5/16   * 144

       =67.5 kw

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin